QUESTION 21 3. The cofunction value of col(50 + x) is: O Cot(40° + x) Ob tan(40 + x) Ootan(40 - x) is: d. none of the abov
Answer #1

3. The cofunction value of Cot (50 +x) = tem {96 – (50 +27] since 1. We know; ento = tan (96-0). ... 2. we can write . tainis ele o = 3.1 x 120 (1: Qmin. s120sec.) = 311x30 = 90T feet. Distance (8) = 2:489017 = 678.5 feet. = 679 feet So, the requir